OpenStudy (anonymous):

Simplify the rational expression: (2x^2+2x-4)/(x^2-5x-14) Please show work! Thanks(:

OpenStudy (anonymous):

2(x-2)/(x-7)

OpenStudy (anonymous):

I'm sorry but i don't understand. Can you please show more work?

OpenStudy (anonymous):

2(x-1)(x+2)/(x-7)(x+2) 2(x-1)/(x-7)

OpenStudy (anonymous):

Thank you sammy!!! :D

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

(2x^2+2x-4)/(x^2-5x-14) = 2(x +2) ( x -1)/ ( x -7) (x +2) = 2 (x-1)/ (x-7)